OpenStudy (anonymous):

what is the square root of 35?

OpenStudy (anonymous):

im just curious. :)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

use your calculator and find out

OpenStudy (anonymous):

i have no idea what a square root but on this movie this girl asked what that was and i was curious soooo.

OpenStudy (nowhereman):

\(\sqrt{35} = \sqrt 5 \cdot \sqrt 7 \approx 5.93\) And the square root of \(x\) is the number \(y\geq 0\) so that \(y^2 = x\)

OpenStudy (anonymous):

jimmy rep its not that easy for me i dont have a calculater and its niot my thig to do math that way i do it in my heaad

OpenStudy (nowhereman):

Which movie, btw?

OpenStudy (anonymous):

well i saw this movie along time ago so i dont know what it is called but i know it was or had something to do with love it was definitly a chick flick and most likly a short of some kind :)

OpenStudy (anonymous):

sorry for any misspelled words :)